suppose E[X]=2 and E[Y]=1. can we calculate E[XY]? explain why or why not?
Answer #1

E[XY] = E(X)*E(Y)

Only if it is given that X any Y are independent of each other.

Then,

E[XY] = 2*1 = 2

But in case they are not independent then the above solution will be incorrect.

Ex:- For not independent event(attached file)oe 2. 6 1 2 12